Olestra is a fat substitute approved by the FDA for use in snack foods. Because there have been anecdotal reports of gastroin===========================================
1 An article describes an experiment in which several types of boxes were compared with respect to compression strength (lb).f=25.09 is wrong.
Answer #1

1)

a)

1st group 2nd group
x=    83 74
n = 500 500
p̂=x/n= 0.1660 0.1480
estimated prop. diff =p̂1-p̂2    = 0.0180
pooled prop p̂ =(x1+x2)/(n1+n2)= 0.1570
std error Se=√(p̂1*(1-p̂1)*(1/n1+1/n2) = 0.0230
test stat z=(p̂1-p̂2)/Se = 0.78
P value   = 0.4354 (from excel:2*normsdist(-0.78)

b)

first sample proportion p1 = 0.2
second sample proportion p2= 0.25
0.05 level critical Z= 1.96
0.1 level critical Zβ= 1.28
n=((zα(√(p1+p2)*(q1+q2)/2)+zβ(√p2q2+p2q2))/(p1-p2))2= 1465


2)Applying ANOVA:

Source of Variation SS df MS F P-value
Between Groups 22781.2546 3 7593.7515 4.4882 0.0145
Within Groups 33838.9750 20 1691.9488
Total 56620.2296 23

f =4.49